A chopper drops boxes of relief goods to a group of people stranded in an island. The goods were released 20.0 m above the sea and landed 40.0 m from a point exactly below where the goods were released. What was the velocity of the goods when they were released? What was the velocity of the chopper?




1
Expert's answer
2021-11-25T16:23:22-0500

"v=\sqrt{2gh}=20~\frac ms,"

"t=\sqrt{\frac{2h}g}=2~s;"

"u=\frac st=20~\frac ms."